Ben walks 5 kilometer to his house from quiapo to buy a new bike return home. He averaged 10 kilometer faster on his bike than on foot. If this total trip took 1hour and 20minutes what is the speed walking in kph hint v=d/t
1
Expert's answer
2021-09-12T19:01:10-0400
"v=\frac{d}{t}\\v=\frac{10}{20}60=30\frac{km}{h}"